OpenStudy (demonchild99):

HELP...i have to make a 8o or im going o be put back in regular school.

OpenStudy (demonchild99):

Given a soda can with a volume of 15 and a diameter of 2, what is the volume of a cone that fits perfectly inside the soda can? (Hint: only enter numerals in the answer blank). Numerical Answers Expected! Answer for Blank 1: Question 2 (Fill-In-The-Blank Worth 4 points) (08.01 MC) The stack on the left is made up of 15 pennies, and the stack on the right is also made up of 15 pennies. If the volume of the stack of pennies on the left is 390 mm3, what is the volume of the stack of pennies on the right in cubic millimeters? (Hint: only enter numerals in the answer blank) Two stacks of pennies are shown. The stack of pennies on the left form a right cylinder while the stack of pennies on the right form a free form cylinder like shape. Numerical Answers Expected! Answer for Blank 1: Question 3 (Multiple Choice Worth 4 points) (08.01 MC) Given a cube with a volume of 36 cm3, what is the volume of a square pyramid that can fit perfectly inside the cube? 3cm3 6cm3 9cm3 12cm3 Question 4 (Multiple Choice Worth 4 points) (08.01 LC) Find the circumference and area of a circle with a diameter of 10 inches. Leave your answers in terms of pi. C = 5π; A = 20π C = 10π; A = 20π C = 10π; A = 25π C = 5π; A = 25π Question 5 (Fill-In-The-Blank Worth 4 points) (08.01 MC) Find the volume of a cylinder with a diameter of 6 inches and a height that is three times the radius. Use 3.14 for pi and round your answer to the nearest hundredth. (Hint: You may only enter numerals, decimal points, and negative signs in the answer blank) Answer for Blank 1:

OpenStudy (jordanloveangel):

ill help u in a sec k

OpenStudy (demonchild99):

ok!

OpenStudy (demonchild99):

this is for the pennies

OpenStudy (jordanloveangel):

Given a cube with a volume of 36 cm3, what is the volume of a square pyramid that can fit perfectly inside the cube? 3cm3 6cm3 9cm3 12cm3 volume of cube = a^3 Volume of Square Pyramid = (a^2 h)/3 the height and base of square pyramid is equal to the side of cube..

OpenStudy (jordanloveangel):

Find the volume of a cylinder with a diameter of 6 inches and a height that is three times the radius. Use 3.14 for pi and round your answer to the nearest hundredth. (Hint: You may only enter numerals, decimal points, and negative signs in the answer blank) Volume of a cylinder = (pi)(r^2)h Diameter = 6. Radius is half the diameter. So radius r = 3 Height is three times the radius. So height h = 3 x r = 3 x 3 = 9 Just plug it into the formula for volume. I got 169.56 Volume = (pi)(r^2)h = (3.14)(3^2)(9) = 3.14 x 9 x 9 = 254.34

OpenStudy (jordanloveangel):

Find the circumference and area of a circle with a diameter of 10 inches. Leave your answers in terms of pi. C = 5π; A = 20π C = 10π; A = 20π C = 10π; A = 25π C = 5π; A = 25π 10π; A = 25π

OpenStudy (demonchild99):

i hate those stupid question marks

OpenStudy (jordanloveangel):

lol refresh ur page

OpenStudy (demonchild99):

kk

OpenStudy (jordanloveangel):

Given a soda can with a volume of 15 and a diameter of 2, what is the volume of a cone that fits perfectly inside the soda can? (Hint: only enter numerals in the answer blank). volume of cylinder=πr^2h Volume of cone=1/3πr^2h=1/3∗15=5 Because r and h are same.

OpenStudy (jordanloveangel):

okay this is all i now

OpenStudy (demonchild99):

so..the answer would be 5? @jordanloveangel

OpenStudy (jordanloveangel):

for which one

OpenStudy (demonchild99):

the last one u did

OpenStudy (jordanloveangel):

volume = PI * r² * height 15 / PI * 1² = height height = 15 / PI = 4.77 Now you need the volume of a cone with height of 4.77 and radius of one. cone volume = (π • r² • h) / 3 cone volume = (3.14159 * 1 * 4.77) / 3 cone volume = 14.9853969576 / 3 cone volume = 4.995 or 5 rounded Here's a cone calculator to check your work.

OpenStudy (jordanloveangel):

http://www.1728.org/volcone.htm

OpenStudy (jordanloveangel):

Here is a quick solution I just noticed: cylinder volume = PI * r² * height cone volume = (PI * r² * height) / 3 The exact same numbers go in the exact same place so we can see that a cone of equal height and radius of a cylinder will have 3 times less volume due to the division by 3 in the cone formula.

OpenStudy (demonchild99):

thank you so much....